Pat Kinch used a racing cycle to travel 20.0 m/s. Suppose Kinch moved at this speed around a circular track. If the combined mass of Kinch and the cycle was 92.0 kg and the average centripetal force was 920 N, what was the radius in meters of the track? (Centripetal Force: Fc = m(v^2)/r ) (Write the number answer only.)



1
Expert's answer
2021-02-25T11:25:36-0500
"F_c=\dfrac{mv^2}{r},""r=\dfrac{mv^2}{F}=\dfrac{92\ kg\cdot(20\ \dfrac{m}{s})^2}{920\ N}=40\ m."